OpenStudy (anonymous):

what is -1/10 - (-2/10)?

pooja195 (pooja195):

2 negative signs next to each other make a postive denominator always remains the same so: \[\huge~\rm~\frac{ -1+2 }{ 10}=?\]

OpenStudy (anonymous):

-1/10?

pooja195 (pooja195):

-1+2=?

OpenStudy (anonymous):

1?

pooja195 (pooja195):

Yes so our fraction would be...?

OpenStudy (anonymous):

1/10

pooja195 (pooja195):

Perfect good job :-)

OpenStudy (anonymous):

ty

pooja195 (pooja195):

yw ^-^